OpenStudy (anonymous):

The volume of a cube is found using the formula L^3 where L is the side length. What is the volume of a cube in cubic feet, that is 3/5 of a foot long?

OpenStudy (anonymous):

27/125 cubic feet? maybe?

OpenStudy (anonymous):

Thats one of the options :D thank you!

OpenStudy (rogue):

\[V_{cube} = s^3, s = 3/5 ft, V = (\frac {3}{5})^3 = \frac {3^3}{5^3} = \frac {27}{125} ft^3\]